Shania would like to give $5 gift cards and $4 teddy bears as party favors. 16 people were invited. Shania has $100 to spend on party favors. Write an inequality to find the number of gift cards x and teddy bears y Shania could purchase.
Answer by richwmiller(17219)

x+y=16
5x+4y<=100
we need only 16 total maximum
5*16=80
4*16=64
so the cost could range from 64 to 80 dollars
so she could have from 0 to 16 teddy bears and 0 to 16 gift cards
x y
16 0
15 1
14 2
13 3
12 4
etc
or
x y
0 16
1 15
2 14
etc